# A vectorial Darboux transformation for integrable matrix versions of the Fokas-Lenells equation

### v1

@@ 6.2 The generic case $\alpha_{1}\neq\alpha_{2}$
$\chi_{1}$ $=$ $\sum_{\kappa}e^{R_{\kappa}x+S_{\kappa}t}c_{\kappa}\,$,
$\chi_{2}$ $=$ $\frac{\|A\|^{2}}{(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}}\Big{(}\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\Gamma^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}I_{n}\Big{)}^{-1}\sum_{\kappa}\Big{(}R_{\kappa}^{2}-\frac{\mathrm{i}\,\,A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,R_{\kappa}-\frac{1}{4}\Gamma^{-2}$
@@
and (5.7), which is (6.22). This “degenerate” case of the Lyapunov equation includes the important class of dark solitons.
6.2.1 The case $n=1$
@@
where $\gamma\neq\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,A^{\dagger}\boldsymbol{\alpha}A/(\alpha_{1}\alpha_{2}\|A\|^{2})$, $\{r_{\kappa}\}$ is a set of different roots of the cubic equation
@@
$s_{\kappa}=\frac{1}{\alpha_{1}\alpha_{2}}\Big{(}\gamma\,r_{\kappa}^{2}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})\gamma\,r_{\kappa}-\frac{1}{2}\alpha_{1}\alpha_{2}\gamma-\frac{1}{4}\gamma^{-1}+\mathrm{i}\,(\alpha_{1}+\alpha_{2})(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2})\Big{)}\,.$
If there are three different roots $r_{\kappa}$, $\kappa=1,2,3$, then the above solution of the linear system is the general solution. Otherwise, the missing solutions still have to be found.
@@
where $r$ is any root of the above cubic equation (6.50) (and $s$ the corresponding $s_{\kappa}$). From (6.50) and its complex conjugate, we get
@@
$\Omega$ $=$ $-\frac{|c|^{2}\,\delta_{1}}{k^{2}}\Big{(}1-\frac{\delta_{1}^{2}\,(4\,\delta_{2}\,\mathrm{Im}(r)+\alpha_{1}\alpha_{2})}{(2\,\delta_{1}\delta_{2}+k\,\alpha_{1}\alpha_{2})^{2}}\Big{)}\,\Big{(}x+\frac{2\,k\,\delta_{2}}{\alpha_{1}\alpha_{2}}\,t+\tilde{c}\Big{)}\,$,
@@
Figure 9: Plots of the absolute values of the two components of a solution from the class in Section 6.2.1 (2), showing a bright-dark soliton. Here we chose $\gamma=-\mathrm{i}\,$ (i.e., $k=1$), $r=-1+\mathrm{i}\,/3$, $\alpha_{1}=-\alpha_{2}=1$, $c=1$, $\tilde{c}=-1$.
6.2.2 The triple root case and rogue waves
@@
The solution of the linear equation for $\chi_{1}$, obtained in Section 6 in terms of only exponential functions, is not complete in case of a multiple root of the cubic equation (6.50), because there are then additional solutions involving polynomials of the independent variables. In this section, we address the special case of a triple root (also see [10, 13]). This leads to quasi-rational solutions of the two-component vector FL equation, which turn out to describe in space and time (formally regarding $t$ as time and $x$ as space coordinate) localized perturbations of the plane wave background.
@@
$\theta=r\,x+\frac{1}{\alpha_{1}\alpha_{2}}\Big{(}r^{2}\gamma-\mathrm{i}\,\,r\gamma(\alpha_{1}+\alpha_{2})-\big{(}\frac{1}{2}\alpha_{1}\alpha_{2}\gamma+\frac{1}{4}\gamma^{-1}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})(A^{\dagger}\alpha A-\frac{1}{2})\big{)}\Big{)}\,t\,$,
@@
It is quasi-rational and includes counterparts of the Peregrine breather solution of the NLS equation, modeling a rogue wave that is localized in space and time.
If $c_{1}=c_{2}=0$, then $u_{1}^{\prime}$ and $u_{2}^{\prime}$ reduce to the plane wave background solutions. Therefore we shall now assume $c_{1}c_{2}\neq 0$. If $c_{1}=0$ and $c_{2}\neq 0$, so that $f_{1}=0$ and $f_{2}=c_{2}\neq 0$, then $\chi_{1}$, $\chi_{2}$ and $\Xi$ only involve polynomials (in the independent variables) that are linear. An example of the corresponding rogue wave is displayed in Fig. 10. If $c_{1}\neq 0$, the polynomials, appearing in $\chi_{1}$, $\chi_{2}$ and $\Xi$, are quadratic. In this case, the components $u_{1}^{\prime}$ and $u_{2}^{\prime}$ form a kind of (elementary) rogue wave pairs. An example is shown in Fig. 11.

### v2

@@ 6.2 The generic plane wave case $\alpha_{1}\neq\alpha_{2}$
$\chi_{1}$ $=$ $\sum_{\kappa}e^{R_{\kappa}x+S_{\kappa}t}c_{\kappa}\,$, (6.48)
$\chi_{2}$ $=$ $\frac{\|A\|^{2}}{(\alpha_{2}-\alpha_{1})\,A_{1}A_{2}}\Big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,\Gamma^{-1}-\frac{\alpha_{1}\alpha_{2}\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}I_{n}\Big)^{-1}\sum_{\kappa}\Big(R_{\kappa}^{2}-\frac{\mathrm{i}\,A^{\dagger}\boldsymbol{\alpha}^{2}A}{A^{\dagger}\boldsymbol{\alpha}A}\,R_{\kappa}-\frac{1}{4}\Gamma^{-2}$ (6.49)
@@
and (5.7), which is (6.22).
6.2.1 $n=1$
@@
assuming $\gamma\neq\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,A^{\dagger}\boldsymbol{\alpha}A/(\alpha_{1}\alpha_{2}\|A\|^{2})$. Here $\{r_{\kappa}\}$ is a set of different roots of the cubic characteristic equation
@@
$s_{\kappa}=\frac{1}{\alpha_{1}\alpha_{2}}\Big(\gamma\,r_{\kappa}^{2}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})\gamma\,r_{\kappa}-\frac{1}{2}\alpha_{1}\alpha_{2}\gamma-\frac{1}{4}\gamma^{-1}+\mathrm{i}\,(\alpha_{1}+\alpha_{2})(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2})\Big)\,.$ (6.63)
If there are three different roots $r_{\kappa}$, $\kappa=1,2,3$, then the above solution of the linear system is the general solution. Otherwise, the missing solutions still have to be found, which is addressed in subsections 6.2.3 and 6.2.4.
@@
where $r$ is any root of the characteristic equation (6.62) (and $s$ the corresponding $s_{\kappa}$). From (6.62) and its complex conjugate, we get
@@
[…]
@@
Figure 11: Plots of the absolute values of the two components of a solution from the class in Section 6.2.1 (ii), showing a bright-dark soliton. Here we chose $\gamma=-\mathrm{i}$ (i.e., $k=1$), $r=-1+\mathrm{i}/3$, $\alpha_{1}=-\alpha_{2}=1$, $c=1$, $\tilde{c}=-1$.
6.2.2 $n=2$ and generic roots
Given two $n=1$ solutions from Section 6.2.1, with spectral parameters $\gamma_{2}\neq-\gamma_{1}^{\ast}$, a corresponding superposition, which also solves the vector FL equation, is easily computed. With the correspondingly generalized restriction, an arbitrary number of $n=1$ solutions can easily be superposed. If, however, the (relative part of the) spectrum condition for the Lyapunov equation does not hold, it is more involved to compute the superposition of $n=1$ solutions. We elaborate an $n=2$ case.
Example 6.13.
Let $\Gamma=\mathrm{diag}(\gamma,-\gamma^{\ast})$ with $\mathrm{Re}(\gamma)\neq 0$. The Lyapunov equation determines the diagonal entries of $\Omega$,
$\Omega_{11}$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}\,\gamma^{\ast}|\Xi_{1}|^{2}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}+|\chi_{11}|^{2}+|\chi_{21}|^{2}\Big)\,$,
$\Omega_{22}$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}\,\gamma|\Xi_{2}|^{2}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}-|\chi_{12}|^{2}-|\chi_{22}|^{2}\Big)\,$,
where
$\Xi=\|A\|\,\Big(\chi_{1x}-\big(\frac{1}{2}\Gamma^{-1}+\mathrm{i}\,\frac{A^{\dagger}\boldsymbol{\alpha}A}{\|A\|^{2}}\,I_{n}\big)\,\chi_{1}\Big)-\mathrm{i}\,(\alpha_{2}-\alpha_{1})\,\frac{A_{1}A_{2}}{\|A\|}\,\chi_{2}\,$,
and the constraints
$\Big(\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Xi\Xi^{\dagger}\,\Gamma^{\dagger}+\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\Big)_{ij}=0\qquad\mbox{for}\quad i\neq j\,.$
We observe that if $r$ is a solution of (6.62), for a given parameter $\gamma$, then $r^{\ast}$ is a solution of (6.62) with $\gamma$ replaced by $\gamma^{\ast}$. Assuming that the three roots of (6.62) are all different [^8] If there is a multiple root, we still obtain exact solutions by restricting the sums to a set of different roots., the linear system is completely solved by (6.48) and (6.49), where now $R_{\kappa}=\mathrm{diag}(r_{\kappa},-r_{\kappa}^{\ast})$, $S_{\kappa}=\mathrm{diag}(s_{\kappa},-s_{\kappa}^{\ast})$, with $s_{\kappa}$ given by (6.63), and $c_{\kappa}=(c_{1\kappa},c_{2\kappa})^{T}$. Then the above constraints boil down to
$\sum_{\kappa=1}^{3}c_{1\kappa}\,c_{2\kappa}^{\ast}\,\big(1+d_{\kappa}^{2}+\mathrm{i}\,\gamma h_{\kappa}^{2}\big)=0\,$,
where
$d_{\kappa}$ $=$ $\frac{1}{(\alpha_{2}-\alpha_{1})A_{1}A_{2}\big(\mathrm{i}\,(1-A^{\dagger}\boldsymbol{\alpha}A)\,A^{\dagger}\boldsymbol{\alpha}A-\gamma\,\alpha_{1}\alpha_{2}\|A\|^{2}\big)}\Big(\gamma\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A\,r_{\kappa}^{2}-\mathrm{i}\,\gamma\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}^{2}A\,r_{\kappa}$
$-\frac{1}{4}\gamma^{-1}\|A\|^{2}A^{\dagger}\boldsymbol{\alpha}A+\frac{1}{2}\mathrm{i}\,(A^{\dagger}\boldsymbol{\alpha}A)^{2}(2A^{\dagger}\boldsymbol{\alpha}A-1)+\frac{1}{2}\mathrm{i}\,(\alpha_{2}-\alpha_{1})^{2}|A_{1}|^{2}|A_{2}|^{2}\Big)\,$,
$h_{\kappa}$ $=$ $\frac{1}{2\|A\|A^{\dagger}\boldsymbol{\alpha}A}\Big(2\,(\|A\|^{2}r_{\kappa}-\mathrm{i}\,k)-\gamma^{-1}\|A\|^{2}-2\mathrm{i}\,d_{\kappa}A_{1}A_{2}\,(\alpha_{2}-\alpha_{1})\Big)\,.$
The differential equations for $\Omega$ read
$\Omega_{12x}$ $=$ $\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Xi_{1x}\Xi_{2}^{\ast}-\frac{1}{2}\gamma^{-2}\chi_{11}\chi_{12}^{\ast}+\gamma^{-1}\Big(\frac{A_{1}^{\ast}A_{2}^{\ast}}{A^{\dagger}\boldsymbol{\alpha}A}(\alpha_{2}-\alpha_{1})\,(\chi_{11x}-\frac{1}{2}\gamma^{-1}\chi_{11})-\chi_{21x}$
$+\frac{\mathrm{i}\,\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\alpha_{1}\alpha_{2}\chi_{21}\Big)\,\chi_{22}^{\ast}\,$,
$\Omega_{12t}$ $=$ $\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\big(\gamma\,\Xi_{2}^{\ast}-\|A\|A^{\dagger}\boldsymbol{\alpha}A\,\chi_{12}^{\ast}\big)\,\Xi_{1}\,$,
$\Omega_{21x}$ $=$ $\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\Xi_{1}^{\ast}\Xi_{2x}-\frac{1}{2}\gamma^{\ast-2}\chi_{11}^{\ast}\chi_{12}-\gamma^{\ast-1}\Big(\frac{A_{1}^{\ast}A_{2}^{\ast}}{A^{\dagger}\boldsymbol{\alpha}A}(\alpha_{2}-\alpha_{1})\,(\chi_{12x}+\frac{1}{2}\gamma^{\ast-1}\chi_{12})-\chi_{22x}$
$+\frac{\mathrm{i}\,\|A\|^{2}}{A^{\dagger}\boldsymbol{\alpha}A}\alpha_{1}\alpha_{2}\chi_{22}\Big)\,\chi_{21}^{\ast}\,$,
$\Omega_{21t}$ $=$ $-\frac{\mathrm{i}}{(A^{\dagger}\boldsymbol{\alpha}A)^{2}}\big(\gamma^{\ast}\Xi_{1}^{\ast}+\|A\|A^{\dagger}\boldsymbol{\alpha}A\chi_{11}^{\ast}\big)\,\Xi_{2}\,$,
which leads to
$\Omega_{12}$ $=$ $-\frac{\mathrm{i}}{s_{1}-s_{2}}\big(c_{11}c_{22}^{\ast}h_{1}(\gamma h_{2}+\|A\|)\,e^{(r_{1}-r_{2})x+(s_{1}-s_{2})t}-c_{12}c_{21}^{\ast}h_{2}(\gamma h_{1}+\|A\|)\,e^{-(r_{1}-r_{2})x-(s_{1}-s_{2})t}\big)$
$-\frac{\mathrm{i}}{s_{1}-s_{3}}\big(c_{11}c_{23}^{\ast}h_{1}(\gamma h_{3}+\|A\|)\,e^{(r_{1}-r_{3})x+(s_{1}-s_{3})t}-c_{13}c_{21}^{\ast}h_{3}(\gamma h_{1}+\|A\|)\,e^{-(r_{1}-r_{3})x-(s_{1}-s_{3})t}\big)$
$-\frac{\mathrm{i}}{s_{2}-s_{3}}\big(c_{12}c_{23}^{\ast}h_{2}(\gamma h_{3}+\|A\|)\,e^{(r_{2}-r_{3})x+(s_{2}-s_{3})t}-c_{13}c_{22}^{\ast}h_{3}(\gamma h_{2}+\|A\|)\,e^{-(r_{2}-r_{3})x-(s_{2}-s_{3})t}\big)$
$-\mathrm{i}\,\sum_{\kappa=1}^{3}c_{1\kappa}c_{2\kappa}^{\ast}h_{\kappa}(\gamma h_{\kappa}+\|A\|)\,t+\frac{1}{2\gamma^{2}A^{\dagger}\boldsymbol{\alpha}A}\Big((\alpha_{2}-\alpha_{1})A_{1}^{\ast}A_{2}^{\ast}\sum_{\kappa=1}^{3}c_{1\kappa}c_{2\kappa}^{\ast}d_{\kappa}(2\gamma r_{\kappa}-1)$
$-\sum_{\kappa=1}^{3}\big(A^{\dagger}\boldsymbol{\alpha}A\,c_{1\kappa}c_{2\kappa}^{\ast}(2\mathrm{i}\,\gamma^{2}r_{\kappa}h_{\kappa}^{2}+2\gamma r_{\kappa}d_{\kappa}^{2}+1)-2\mathrm{i}\,\gamma\alpha_{1}\alpha_{2}\|A\|^{2}c_{1\kappa}c_{2\kappa}^{\ast}d_{\kappa}^{2}\big)\Big)\,x+\tilde{c}_{12}\,$,
$\Omega_{21}$ $=$ $\frac{\mathrm{i}}{s_{1}^{\ast}-s_{2}^{\ast}}\big(c_{11}^{\ast}c_{22}h_{2}^{\ast}(\gamma^{\ast}h_{1}^{\ast}+\|A\|)\,e^{(r_{1}^{\ast}-r_{2}^{\ast})x+(s_{1}^{\ast}-s_{2}^{\ast})t}-c_{12}^{\ast}c_{21}h_{1}^{\ast}(\gamma^{\ast}h_{2}^{\ast}+\|A\|)\,e^{-(r_{1}^{\ast}-r_{2}^{\ast})x-(s_{1}^{\ast}-s_{2}^{\ast})t}\big)$
$+\frac{\mathrm{i}}{s_{1}^{\ast}-s_{3}^{\ast}}\big(c_{11}^{\ast}c_{23}h_{3}^{\ast}(\gamma^{\ast}h_{1}^{\ast}+\|A\|)e^{(r_{1}^{\ast}-r_{3}^{\ast})x+(s_{1}^{\ast}-s_{3}^{\ast})t}-c_{13}^{\ast}c_{21}h_{1}^{\ast}(\gamma^{\ast}h_{3}^{\ast}+\|A\|)\,e^{-(r_{1}^{\ast}-r_{3}^{\ast})x-(s_{1}^{\ast}-s_{3}^{\ast})t}\big)$
$+\frac{\mathrm{i}}{s_{2}^{\ast}-s_{3}^{\ast}}\big(c_{12}^{\ast}c_{23}h_{3}^{\ast}(\gamma^{\ast}h_{2}^{\ast}+\|A\|)\,e^{(r_{2}^{\ast}-r_{3}^{\ast})x+(s_{2}^{\ast}-s_{3}^{\ast})t}-c_{13}^{\ast}c_{22}h_{2}^{\ast}(\gamma^{\ast}h_{3}^{\ast}+\|A\|)\,e^{-(r_{2}^{\ast}-r_{3}^{\ast})x-(s_{2}^{\ast}-s_{3}^{\ast})t}\big)$
$+\mathrm{i}\,\sum_{\kappa=1}^{3}c_{1\kappa}^{\ast}c_{2\kappa}h_{\kappa}^{\ast}(\gamma^{\ast}h_{\kappa}^{\ast}+\|A\|)\,t+\frac{1}{2\gamma^{\ast 2}A^{\dagger}\boldsymbol{\alpha}A}\Big((\alpha_{2}-\alpha_{1})A_{1}^{\ast}A_{2}^{\ast}\sum_{\kappa=1}^{3}c_{1\kappa}^{\ast}c_{2\kappa}d_{\kappa}^{\ast}(2\gamma^{\ast}r_{\kappa}^{\ast}-1)$
$+\sum_{\kappa=1}^{3}\big(A^{\dagger}\boldsymbol{\alpha}A\,c_{1\kappa}^{\ast}c_{2\kappa}(2\mathrm{i}\,\gamma^{\ast 2}r_{\kappa}^{\ast}h_{\kappa}^{\ast 2}-2\gamma^{\ast}r_{\kappa}^{\ast}d_{\kappa}^{\ast 2}-1)-2\mathrm{i}\,\gamma^{\ast}\alpha_{1}\alpha_{2}\|A\|^{2}c_{1\kappa}^{\ast}c_{2\kappa}d_{\kappa}^{\ast 2}\big)\Big)\,x+\tilde{c}_{21}\,$,
with complex constants $\tilde{c}_{12}$ and $\tilde{c}_{21}$. The remaining condition (6.22) becomes
$\gamma\,(\tilde{c}_{12}-\tilde{c}_{21}^{\ast})=\sum_{\kappa=1}^{3}\big(c_{1\kappa}\,c_{2\kappa}^{\ast}+d_{1\kappa}\,d_{2\kappa}^{\ast}\big)\,.$
A class of solutions of the two-component FL equation is now obtained by inserting the expressions for $\chi_{1}$, $\chi_{2}$, $\Xi$ and $\Omega$ in (6.21). In addition to the exponential dependence on the independent variables via $e^{\pm(r_{i}-r_{j})x\pm(s_{i}-s_{j})t}$ and their complex conjugates, the solution also exhibits a rational dependence, a typical feature in those cases, where the spectrum condition for the Lyapunov equation does not hold. $\Box$
Example 6.14.
Let $\Gamma$ be the $2\times 2$ lower triangular Jordan block with eigenvalue $\gamma$. Then we have
$R_{\kappa}=\left(\begin{array}[]{cc}r_{\kappa}&0\\ \tilde{r}_{\kappa}&r_{\kappa}\end{array}\right)\,,\qquad S_{\kappa}=\left(\begin{array}[]{cc}s_{\kappa}&0\\ \tilde{s}_{\kappa}&s_{\kappa}\end{array}\right)\,$,
with
$\tilde{r}_{\kappa}$ $=$ $\frac{8\mathrm{i}\,\gamma A^{\dagger}\boldsymbol{\alpha}^{2}A\,(\gamma r_{\kappa}-1)+4\gamma^{2}\alpha_{1}\alpha_{2}\,(2A^{\dagger}\boldsymbol{\alpha}A-1)+4\mathrm{i}\,\gamma\,(\alpha_{1}+\alpha_{2})-4\gamma r_{\kappa}\,(1+\gamma r_{\kappa})+3}{2\gamma^{2}\big(4\mathrm{i}\,\gamma A^{\dagger}\boldsymbol{\alpha}^{2}A-8\mathrm{i}\,\gamma^{2}r_{\kappa}\,(\alpha_{1}+\alpha_{2})-4\gamma^{2}\alpha_{1}\alpha_{2}+12\gamma^{2}r_{\kappa}^{2}-4\gamma r_{\kappa}-1\big)}\,$,
$\tilde{s}_{\kappa}$ $=$ $-\frac{1}{4\alpha_{1}\alpha_{2}\gamma^{2}}\Big(4\mathrm{i}\,\gamma^{2}(r_{\kappa}+\gamma\tilde{r}_{\kappa})\,(\alpha_{1}+\alpha_{2})+2\gamma^{2}\alpha_{1}\alpha_{2}-8\gamma^{3}r_{\kappa}\tilde{r}_{\kappa}-4\gamma^{2}r_{\kappa}^{2}-1\Big)\,.$
Assuming that the three roots $r_{\kappa}$, $\kappa=1,2,3$, of (6.62) are all different, the linear system is completely solved by (6.48) and (6.49). Then $\Xi$ is easily computed and it remains to determine $\Omega$, in order to be able to elaborate the resulting solution of the two-component vector FL equation, which is a second order version of that in Section 6.2.1 with the same $\gamma$. If $\mathrm{Re}(\gamma)\neq 0$, then $\Omega$ is fully determined by the Lyapunov equation, see Example A.2. If $\mathrm{Re}(\gamma)=0$, the differential equations for $\Omega$ have to be solved, which is more tedious. We omit the details. $\Box$
6.2.3 The triple root case
@@
The solution of the linear equation for $\chi_{1}$, obtained in Section 6.2 in terms of only exponential functions, is not complete in case of a multiple root of the characteristic equation (6.62), because there are then additional solutions involving polynomials of the independent variables. In this section, we address the special case of a triple root (also see [11, 16]). This leads to genuine [^9] In contrast to solutions that are equivalent to solutions of the scalar FL equation. quasi-rational solutions of the two-component vector FL equation, which turn out to describe in space and time (formally regarding $t$ as time and $x$ as space coordinate) localized perturbations of the plane wave background.
@@
$\theta=r\,x+\frac{1}{\alpha_{1}\alpha_{2}}\Big(r^{2}\gamma-\mathrm{i}\,r\gamma(\alpha_{1}+\alpha_{2})-\big(\frac{1}{2}\alpha_{1}\alpha_{2}\gamma+\frac{1}{4}\gamma^{-1}-\mathrm{i}\,(\alpha_{1}+\alpha_{2})(A^{\dagger}\boldsymbol{\alpha}A-\frac{1}{2})\big)\Big)\,t\,$, (6.71)
@@
We note that $e^{\theta}$ drops out in the expression for $u^{\prime}$, so that the solution $u^{\prime}$ is quasi-rational.
If $c_{1}=c_{2}=0$, then $u_{1}^{\prime}$ and $u_{2}^{\prime}$ reduce to the plane wave background solutions.
If $c_{1}=0$ and $c_{2}\neq 0$, so that $f_{1}=0$ and $f_{2}=c_{2}\neq 0$, then $\chi_{1}$, $\chi_{2}$ and $\Xi$ only involve polynomials (in the independent variables) that are linear. In this case we obtain the single Peregrine-type soliton solution
$u_{\mu}^{\prime}=\frac{1}{2}A_{\mu}\,e^{\mathrm{i}\,\varphi_{\mu}}\Big(-1\pm(-1)^{\mu}\,\mathrm{i}\,\sqrt{3}-\frac{4\,\mathrm{Re}(\gamma)A^{\dagger}\boldsymbol{\alpha}A\,F_{\mu}}{D\,A_{\mu}}\Big)\qquad\quad\mu=1,2\,$,
where
$F_{\mu}$ $=$ $\varrho_{3}\,\Big(A_{\mu}\big(\mathrm{i}\,\gamma^{\ast}\varrho_{2}^{\ast}\varrho_{3}\|A\|^{2}+\varrho_{1}^{\ast}\varrho_{4}(A^{\dagger}\boldsymbol{\alpha}A)^{2}\big)+\tilde{A}_{\mu}\big(\mathrm{i}\,\gamma^{\ast}\varrho_{2}^{\ast}\varrho_{3}\varrho_{4}^{\ast}\|A\|^{2}-\varrho_{1}^{\ast}(\mathrm{i}\,\gamma^{\ast}|\varrho_{3}|^{2}\|A\|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2})\big)\Big)\,\rho$
$+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(A_{\mu}+\varrho_{4}^{\ast}\tilde{A}_{\mu})\,(\varrho_{1}\varrho_{3}\varrho_{4}^{\ast}-\varrho_{2}|\varrho_{4}|^{2}-\varrho_{2})\,\rho^{\ast}+\varrho_{3}A_{\mu}\big(\mathrm{i}\,\gamma^{\ast}|\varrho_{2}|^{2}\|A\|^{2}+|\varrho_{1}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2}\big)$
$-\tilde{A}_{\mu}\big(\varrho_{1}^{\ast}\varrho_{2}(\mathrm{i}\,\gamma^{\ast}|\varrho_{3}|^{2}\|A\|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1+|\varrho_{4}|^{2}))-\varrho_{3}\varrho_{4}^{\ast}(\mathrm{i}\,\gamma^{\ast}|\varrho_{2}|^{2}\|A\|^{2}+|\varrho_{1}|^{2}(A^{\dagger}\boldsymbol{\alpha}A)^{2})\big)\,$,
$D$ $=$ $\big(\mathrm{i}\,\gamma^{\ast}|\varrho_{3}|^{2}\|A\|^{2}+(A^{\dagger}\boldsymbol{\alpha}A)^{2}(1+|\varrho_{4}|^{2})\big)\,\Big(\big((A^{\dagger}\boldsymbol{\alpha}A)^{2}(1+|\varrho_{4}|^{2})+\mathrm{i}\,\gamma^{\ast}|\varrho_{3}|^{2}\|A\|^{2}\big)\,|\rho|^{2}$
$+\big((A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{1}^{\ast}\varrho_{4}+\mathrm{i}\,\gamma^{\ast}\varrho_{2}^{\ast}\varrho_{3}\|A\|^{2}\big)\,\rho+\big((A^{\dagger}\boldsymbol{\alpha}A)^{2}\varrho_{1}\varrho_{4}^{\ast}+\mathrm{i}\,\gamma^{\ast}\varrho_{2}\varrho_{3}^{\ast}\|A\|^{2}\big)\,\rho^{\ast}$
$+(A^{\dagger}\boldsymbol{\alpha}A)^{2}|\varrho_{1}|^{2}+\mathrm{i}\,\gamma^{\ast}|\varrho_{2}|^{2}\|A\|^{2}\Big)\,$,
with $\tilde{A}_{\mu}=(-1)^{\mu}A_{3-\mu}^{\ast}$, $\rho=c_{2}\,(x+(3\alpha_{1}\alpha_{2})^{-1}\tilde{\gamma}\,t)+c_{3}$, and
$\varrho_{1}$ $=\pm\frac{4\sqrt{3}\,c_{2}\,(\nu_{1}^{2}-\nu_{2}^{2})}{A_{1}A_{2}\nu_{2}^{2}\,(2\nu_{1}-\nu_{2})^{2}}\,$, $\qquad\varrho_{2}$ $=\frac{c_{2}\,(\nu_{1}^{2}-4\nu_{1}\nu_{2}+\nu_{2}^{2})}{(2\nu_{1}-\nu_{2})^{2}}\,$,
$\varrho_{3}$ $=\mp\frac{(\nu_{1}+\nu_{2})\,(\nu_{1}^{2}-4\nu_{1}\nu_{2}+\nu_{2}^{2})}{6\nu_{1}(2\nu_{1}-\nu_{2})}\,$, $\qquad\varrho_{4}$ $=\frac{2\sqrt{3}(\nu_{1}+\nu_{2})^{2}(\nu_{1}^{2}-\nu_{1}\nu_{2}+\nu_{2}^{2})}{9A_{1}A_{2}\nu_{1}^{2}\nu_{2}^{2}(2\nu_{1}-\nu_{2})}\,.$
Here $\nu_{1}=\sqrt{3}(\alpha_{1}-\alpha_{2})\pm\mathrm{i}\,(\alpha_{1}+\alpha_{2}),\,\nu_{2}=\sqrt{3}(\alpha_{1}-\alpha_{2})\mp\mathrm{i}\,(\alpha_{1}+\alpha_{2})$. The $\pm$ corresponds to that in the denominator of (6.69). This solution has been explored in particular in [10], where it has been shown that it exhibits surprisingly high amplitudes for particular parameter values, in particular as compared with the Manakov (two-component vector NLS) system (reached in the limit $\epsilon\to 0$, see Remark 6.4 for this parameter). Indeed, choosing for example $c_{2}=c_{3}=1$, we find that the absolute values of the components of $u^{\prime}$ at the point
$(x_{0},t_{0})=\Big(-1\mp\frac{\sqrt{3}\,\big((\alpha_{1}-\alpha_{2})^{4}-\alpha_{1}^{2}\alpha_{2}^{2}\big)}{6\alpha_{1}\alpha_{2}(\alpha_{1}^{3}-\alpha_{2}^{3})}\,,\mp\frac{\sqrt{3}\,\big(\alpha_{1}^{2}-\alpha_{1}\alpha_{2}+\alpha_{2}^{2}\big)^{3}}{6\alpha_{1}\alpha_{2}(\alpha_{1}^{3}-\alpha_{2}^{3})}\Big)$
are given by
$|u_{1}^{\prime}|=|A_{1}|\,\frac{|\alpha_{1}-5\alpha_{2}|}{|\alpha_{1}+\alpha_{2}|}\,,\qquad|u_{2}^{\prime}|=|A_{2}|\,\frac{|5\alpha_{1}-\alpha_{2}|}{|\alpha_{1}+\alpha_{2}|}\,.$
Note that (6.70) requires $\alpha_{\mu}>0$. The last equations show that the peak amplitude can approach nearly five times the background level, which is consistent with the results reported in [10]. An example of a rogue wave from this class of solutions is displayed in Fig. 12.